1. In a study of consumer confidence among middle-class families, 450 were interviewed in the metropolitan area of the country. When asked about having to cut back on discretionary purchases of big-ticket items, 180 of 450 responded that their families had done so over the previous 6 months.


A. What is the 90% confidence interval estimate of the population proportion p of this type of household cutting back on discretionary spending?


B. What sample size would you recommend to achieve a slandered error of 0.03?


1
Expert's answer
2021-12-27T17:59:06-0500

Solution:

Given, n=450, X=180

"\hat p=X/n=180/450=0.4
\\ \hat q=1-\hat p=1-0.4=0.6"

(A):

At 90% CI, the z is:

"\alpha=1-90\%=1-0.9=0.1
\\ \alpha/2=0.1/2=0.05
\\ z_{\alpha/2}=z_{0.05}=1.645"

Margin of error = E =

"z_{\alpha/2}\times \sqrt{\hat p \hat q/n}
\\=1.645\times \sqrt{0.4\times0.6/450}
\\=0.0379"

The 90% CI for population proportion is

"=\hat p-E<p<\hat p+E
\\=0.4-0.379<p<0.4+0.379
\\=0.021<p<0.779"

(B):

Standard error of proportion "=\sqrt{\hat p \hat q/n}"

"0.03=\sqrt{0.4\times 0.6/n}
\\ \Rightarrow 0.0009=0.24/n
\\\Rightarrow n=266.666...
\\ \Rightarrow n\approx267"